# Mildly Super-Eddington Accretion Onto Slowly-Spinning Black Holes Explains the X-Ray Weakness of the Little Red Dots

### v1

@@ 4.1 X-ray Bolometric Corrections
Figure 1: Comparison of the $2-10$ keV bolometric correction as a function of the logarithm of the Eddington ratio in the restframe (left panel) and in the observed $2-10$ keV, with the source placed at the median redshift of $z=6$. Different colors indicate varying inclination angles from the pole (degrees), and different markers represent different spin values: triangles for $a=0.9$, circles for $a=0.68$, and stars for $a=0$. Mildly super-Eddington accretion rates ($41.4<\,{f_{\rm Edd}}<$ 4) onto slowly spinning black holes, if observed at large inclination angles from the pole, lead to extremely large values of the bolometric correction, especially if the SMBH is redshifted at $z>0$.
@@
We chose a representative redshift of $z=6$ as the median value computed from the most comprehensive catalogs of LRDs available to date (Kocevski et al. 2024, where the median redshift is $z=6.1$, and Akins et al. 2024, where the median redshift is $z=6.5$). We note that using a value of $z\sim 5$, which is closer to the median redshift of the sources presented in Harikane et al. (2023) and Maiolino et al. (2023), does not change our results in any significant way.
@@
As the LRDs are high-$z$ sources, we are primarily interested in the observed bolometric corrections calculated for a source at the median redshift of $z=6$; they are displayed in the right panel of Fig. 1. While the general trends with Eddington ratio, spin, and inclination angle do not change, the values of the bolometric corrections are substantially larger, indicating an astounding observed X-ray weakness. In particular, $2/3$ of all the SEDs with mildly super-Eddington rates (i.e., $\,{f_{\rm Edd}}<4$) are characterized by $k_{X}^{z=6}>10^{3}$, with some of them even reaching $k_{X}^{z=6}>10^{4}$. These values are compatible even with the most extreme X-ray weak source found to date (Juodžbalis et al., 2024a).
In summary, GRRMHD simulations of super-Eddington accreting SMBHs show that the emitted SEDs are soft in the X-rays, especially for mildly super-Eddington accretion rates ($41.4<\,{f_{\rm Edd}}<$ 4) onto slowly spinning black holes if observed at large inclination angles from the pole. If these SEDs, which are already intrinsically X-ray weak, are observed at $z\sim 6$, they reach staggering values of the observed X-ray bolometric correction, $>2$ orders of magnitude higher than (restframe) reference values for standard AGN.
@@ 4.2 Optical-UV to X-ray Ratios: the $\,{\alpha_{\rm ox}}$
which is characterized by a mean value of $\,{\alpha_{\rm ox}}\approx 1.5$ for standard AGN (Lusso et al., 2010).
Figure 2 displays that most of our SEDs are characterized by values of the $\,{\alpha_{\rm ox}}$ that are significantly different from the standard $\,{\alpha_{\rm ox}}(L_{2500})$ relation presented in Lusso et al. (2010). In particular, $\sim 55\%$ of our SEDs are outside the dispersion region for typical low-$z$ sources. In extreme cases, the (restframe) luminosity at $2$ keV is $\sim 30$ times weaker than expected from a standard AGN with the same optical-UV luminosity. Interestingly, our models show increasing X-ray weakness, instead of X-ray brightening, for increasing UV luminosity.
Figure 2: The relationship between $\,{\alpha_{\rm ox}}$ and $\log_{10}L_{2500}$ for our super-Eddington SEDs. The dashed red line (with its intrinsic scatter) represents the standard $\,{\alpha_{\rm ox}}(L_{2500\,\mathrm{\AA}})$ relation from Lusso et al. (2010). Many of our super-Eddington SEDs are characterized by values of the $\,{\alpha_{\rm ox}}$ that are significantly lower, indicating a severe lack of X-ray emission. Our SEDs become X-ray weaker, instead of brighter, with increasing optical-UV luminosity.
@@ 4.3 Characterizing the X-ray Slope: the Photon Index
The photon index, $\Gamma$, characterizes the spectral slope of the power-law component of the SED, and it is a valuable measure to investigate how soft (or hard) an SED is at high energy. Assuming a power-law component $F_{\nu}\propto\nu^{\alpha}$, where $\alpha$ is the spectral index, the photon index $\Gamma$ is usually defined as $\Gamma=\alpha+1$, which leads to:
@@
Our values of $\Gamma$ resemble the highest among those recently measured in NLSy1 galaxies by eROSITA (Grünwald et al., 2023). Maiolino et al. (2024) correctly suggested that an intrinsic X-ray weakness in the LRDs could be caused by a steep, extremely X-ray soft SED, similar to that of NLSy1. Our study confirms that this is the case. Care needs to be taken when choosing the photon index $\Gamma$ to characterize the (restframe) bolometric corrections of the LRDs. Choosing too low values, especially if $\Gamma<2$, will lead to overestimating the restframe bolometric correction. Values in the range $\Gamma=1.7-1.9$ have been thus far used in the literature to characterize the X-ray weakness of the LRDs (Yue et al., 2024; Maiolino et al., 2024; Ananna et al., 2024).
@@ 4.4 The Cause of X-ray Weakness: Mildly Super-Eddington, Low Spin, and Beaming
Despite a similar Eddington ratio, the SEDs are markedly different. In particular, the X-ray strong SED is characterized by a much harder X-ray spectrum, with a significant bump extending at (restframe) frequencies of $10^{19}-10^{20}$ Hz for viewing angles close to the pole (i.e., $<30^{\circ}$). Even the SED averaged over the solid angle (shown in red) is hard in the X-ray. On the contrary, the X-ray weak SED does not show any high-energy bump, except, minimally, for a viewing angle of $10^{\circ}$. Hence, the emission is much softer in the X-rays.
Figure 4: Top Row: Comparison (at $z=0$) between typical X-ray weak (left, mildly super-Eddington with zero spin) and X-ray strong (right, mildly super-Eddington with high spin) SEDs. Bottom Row: Same as in the top row, but at $z=6$. The 8 gray lines represent the spectrum for lines of sight ranging from $80^{\circ}$ (bottom) to $10^{\circ}$ (top) from the pole. The red line indicates the spectrum averaged over $4\pi$ steradians. In the bottom panel, the arrows indicate typical flux limits for deep-field X-ray surveys with Chandra ($2-10$ keV) and AXIS ($0.2-10$ keV).
@@
The situation is radically different for a source characterized by an X-ray weak SED. Chandra can barely detect only sources observed near the pole, and it is $\sim 2$ orders of magnitude in flux away from detecting the mean SED. At the flux limit of $10^{-17}\,\rm erg\,s^{-1}\,cm^{-2}$, the mean SED is significantly softer than $0.2$ keV, rendering these sources impossible to detect with Chandra.
@@
Next-generation X-ray observatories, such as AXIS (Reynolds et al., 2023; Cappelluti et al., 2024), can detect such weak sources. Assuming an AXIS deep field with a flux limit of $10^{-18}\,\rm erg\,s^{-1}\,cm^{-2}$ (Marchesi et al., 2020; Cappelluti et al., 2024), and an energy range of $0.2-10$ keV, AXIS will detect the mean SED of a typical X-ray weak source at $z\sim 6$ easily and, most likely, from any viewing angle.
@@ 4.5 X-ray Weak and Strong Accretion Structures
Figure 6: Comparison between accretion structures for our reference cases of X-ray weak (left) and X-ray strong (right) SEDs (see Sec. 4.4). The rows are, from top to bottom: temperature, gas density, and $\beta\gamma$. X-ray strong accretion structures are significantly hotter and more relativistic.
@@
In general, for sub-Eddington accretion rates, the disk is thinner (Novikov & Thorne, 1973; Shakura & Sunyaev, 1973), and a more significant fraction of the hot inner region is visible, leading to a more substantial X-ray component in the observed SED. At super-Eddington accretion rates, although the total luminosity increases, the X-ray contribution diminishes because the hot regions are more obscured by the thicker disk. This angle dependence implies that the apparent fraction of AGNs with intrinsic strong X-ray emission may be underestimated, as many may be viewed at angles where X-rays are suppressed. This is consistent with the explanation of the X-ray weakness provided by Maiolino et al. (2024), which requires significant covering factors.
The main differences in our super-Eddington simulations are shown in the temperature field (top row) and the $\beta\gamma$ field (bottom row), where $\beta=v/c$ and $\gamma$ is the Lorentz factor. The X-ray strong accretion structures display significantly higher values for the temperature and $\beta\gamma$, indicating that the jet is relativistic. The combination of these two effects leads to enhanced X-ray emission. Higher spin values generally lead to more efficient jet production (Blandford et al., 2019), as the spin influences the inner accretion disk’s structure and energy extraction efficiency (see Ricarte et al. 2023 for detailed results on jet efficiencies as a function of black hole spin for super-Eddington accretion).
The temperature distribution within the accretion disk is particularly revealing: the hottest regions, with $T>10^{8}\,\rm K$, are located near the SMBH and are the primary sources of X-ray emission. Hence, observers aligned with the jet would predominantly see high-energy X-rays. Conversely, observers at larger inclination angles are screened from the hottest gas by the geometrically thick disk and see cooler regions, resulting in spectra dominated by UV and optical emissions.
@@ 4.6 Explaining the X-ray Weakness of the LRDs
We can now apply our super-Eddington SEDs to the case of the LRDs.
Figure 7: In the plane $\log_{10}L_{\mathrm{bol}}$ vs. X-ray bolometric correction (in the observed $2-10$ keV range, for $z=6$), we compare the X-ray weakness of our super-Eddington SEDs with that of high-$z$ LRD populations, and some well-known low-$z$ sample. Teal dots represent the bolometric corrections derived in this study from super-Eddington SEDs at $z=6$ (i.e., the median redshift from LRDs catalogs). Note that the bolometric luminosities of our SEDs are compatible with those observed. Orange circles indicate LRD populations at $114<z<$ 11 from Maiolino et al. (2024) and Yue et al. (2024), with arrows showing lower limits, while the red star marks the stacked data point from Maiolino et al. (2024), which provides the most robust constraint. Note that each data point has been shifted to the observed $2-10$ keV range to be comparable to our SEDs. The solid black line (with its spread) represents the restframe $k_{X}(L_{\mathrm{bol}})$ relation from Duras et al. (2020). The grey distribution in the background shows low-$z$ Type 1 AGN data from Lusso et al. (2020).
Figure 7 shows the bolometric correction $k_{X}^{z=6}$ in the observed $2-10$ keV frame for our sample of super-Eddington SEDs, calculated at $z=6$. Note that the bolometric luminosities of our SEDs are compatible with those observed. We display the entire sample of $12$ simulations $\times$ $8$ viewing angles, totaling $96$ values of the bolometric correction. These are compared to (i) lower limits on the bolometric corrections for LRDs at $114<z<$ 11 derived from Maiolino et al. (2024) and Yue et al. (2024); (ii) the X-ray stacking of the Maiolino et al. (2024) sample, which provides the most robust constraint; (iii) the standard (restframe) relation between the bolometric luminosity and the $2-10$ keV bolometric correction presented in Duras et al. (2020); (iv) a density distribution summarizing the (restframe) location in the $L_{\mathrm{bol}}-k_{X}$ plane of standard, low-$z$ Type 1 AGN from Lusso et al. (2020). Note that the data points from Maiolino et al. (2024) and Yue et al. (2024) have been shifted to the observed $2-10$ keV range to be comparable to our SEDs, using the redshift of each source, and $\Gamma=1.7$ and $\Gamma=1.8$, respectively, as indicated in their study.
@@
Remarkably, our models predict that the restframe optical and UV spectra would still exhibit BLRs, potentially with higher velocities and implied line widths due to strong winds from the super-Eddington accreting SMBH. There is recent observational evidence of strong winds launched from super-Eddington accreting AGN (Vietri et al., 2020, 2022; Ding et al., 2022), supported by robust investigations on the theoretical side (Yang et al., 2014; Sądowski et al., 2015; Sądowski & Narayan, 2016; Yang et al., 2023; Zhang et al., 2024). This implies that BLRs could appear artificially broadened by these winds, leading to an overestimation of the black hole mass and an underestimation of the Eddington ratio.

### v2

@@ 4.1 X-ray Bolometric Corrections
Figure 1: Comparison of the $2-10$ keV bolometric correction as a function of the logarithm of the Eddington ratio in the restframe (left panel) and in the observed $2-10$ keV band when the source placed at the median redshift of $z=6$ (right panel). Different colors indicate varying inclination angles from the pole (degrees), and different markers represent different spin values: triangles for $a=0.9$, circles for $a=0.68$, and stars for $a=0$. Mildly super-Eddington accretion rates ($41.4<\,{f_{\rm Edd}}<$ 4) onto slowly spinning black holes, if observed at large inclination angles from the pole, lead to extremely large values of the bolometric correction, especially if the SMBH is highly redshifted.
@@
We chose a representative redshift of $z=6$ as the median value computed from the most comprehensive catalogs of LRDs available to date (Kocevski et al. 2024, where the median redshift is $z=6.1$, Akins et al. 2024, where it is $z=6.5$, and Kokorev et al. 2024, where it is $z=5.6$). We note that using a value of $z\sim 5$, which is closer to the median redshift of the sources presented in Harikane et al. (2023) and Maiolino et al. (2023), does not change our results in any significant way.
@@
As the LRDs are high-$z$ sources, we are primarily interested in the observed bolometric corrections calculated for a source at the median redshift of $z=6$; they are displayed in the right panel of Fig. 1. While the general trends with Eddington ratio, spin, and inclination angle do not change, the values of the bolometric corrections are substantially larger, indicating an astounding observed X-ray weakness. In particular, $2/3$ of all the SEDs with mildly super-Eddington rates (i.e., $\,{f_{\rm Edd}}<4$) are characterized by $k_{X}^{z=6}>10^{3}$, with some of them even reaching $k_{X}^{z=6}>10^{4}$. These values can explain even GN-28074 (Juodžbalis et al., 2024a), the most extreme X-ray weak AGN found to date by JWST (see Fig. 7).
The significant variations in the X-ray bolometric corrections (at a fixed redshift) are due to a combination of different factors. Generally, X-rays are produced by hot gas with temperature $T>10^{7}$ K. In the present simulations, such gas is found exclusively in the jet, both at the jet base and along the walls of the funnel. Higher spin values lead to a more prominent jet (Blandford et al., 2019); correspondingly, non-spinning black holes are the most X-ray-quiet. Jet emission is beamed, hence the observed X-ray emission is strongly dependent on the inclination angle. The Eddington ratio determines the properties of the accretion disk, with higher temperatures being associated with higher values of $\,{f_{\rm Edd}}$, which leads to a higher level of X-ray emission. Section 4.5 describes how the structure of the accretion flow determines the observed X-ray emission.
In summary, GRRMHD simulations of super-Eddington accreting SMBHs show that the observed SEDs are soft in the X-rays, especially for mildly super-Eddington accretion ($41.4<\,{f_{\rm Edd}}<$ 4) on slowly spinning black holes when observed at large inclination angles from the pole. If these SEDs, which are already intrinsically X-ray weak, are observed at $z\sim 6$, they reach staggering values of the observed X-ray bolometric correction, $>2$ orders of magnitude higher than (restframe) reference values for standard AGN.
A note on terminology is warranted. In this work, an "intrinsically X-ray weak" system is defined as one where reduced X-ray emission arises from the intrinsic physical properties of the accretion disk, jet, or corona, or due to the influence of structures close to the black hole (within 100 $r_{g}$) on the emitted radiation. This definition excludes the effects of obscuration from structures at larger radii, such as a torus. Thus, we consider the system within 100 $r_{g}$ as a whole, where few high-energy photons are produced, leading to an intrinsically weak X-ray spectrum. Relativistic beaming (see Sec. 4.4) is not included in this definition; while beaming can enhance the observed X-ray flux from a weak system, it cannot fundamentally alter the intrinsic emission to produce an X-ray strong system.
We conclude by pointing out that the behavior of super-Eddington accretion in the SANE regime is generally similar to that of MAD systems with zero spin, both leading to X-ray-weak outcomes. This is consistent with the results from Curd & Narayan (2019), where both SANE ($a=0$) and MAD ($a=0$) models exhibit X-ray weakness, while only the MAD ($a=0.9$) model is X-ray strong. Therefore, the MAD state is not critical for explaining the X-ray weakness observed in the LRDs, and a SANE accretion state would yield similar results. In the MAD case, a low spin must be invoked to reduce X-ray luminosity, but SANE systems do not require this additional constraint.
@@ 4.2 Optical-UV to X-ray Ratios: the $\,{\alpha_{\rm ox}}$
which is characterized by a mean value of $\,{\alpha_{\rm ox}}\approx 1.5$ for standard high-luminosity AGN (Lusso et al., 2010).
Figure 2 displays that most of our SEDs are characterized by values of the $\,{\alpha_{\rm ox}}$ that are significantly different from the standard $\,{\alpha_{\rm ox}}(L_{2500})$ relation presented in Lusso et al. (2010). In particular, $\sim 55\%$ of our SEDs are outside the dispersion region for typical low-$z$ sources. In extreme cases, the (restframe) luminosity at $2$ keV is $\sim 20$ times weaker than expected from a standard AGN with the same optical-UV luminosity. Interestingly, our models show increasing X-ray emission, instead of X-ray weakening, for increasing UV luminosity.
Figure 2: The relationship between $\,{\alpha_{\rm ox}}$ and $\log_{10}L_{2500}$ for our super-Eddington SEDs. The dashed red line (with its intrinsic scatter) represents the standard $\,{\alpha_{\rm ox}}(L_{2500\,\mathrm{\AA}})$ relation from Lusso et al. (2010). Many of our super-Eddington SEDs are characterized by values of the $\,{\alpha_{\rm ox}}$ that are significantly higher, indicating a severe lack of X-ray emission. Our SEDs become X-ray stronger, instead of weaker, with increasing optical-UV luminosity.
@@ 4.3 Characterizing the X-ray Slope: the Photon Index
The photon index, $\Gamma$, characterizes the spectral slope of the power-law component of the SED, and it is a valuable measure to investigate how soft (or hard) an SED is at high energy. Assuming a power-law component $F_{\nu}\propto\nu^{-\alpha}$, where $\alpha$ is the spectral index, the photon index $\Gamma$ is usually defined as $\Gamma=\alpha+1$, which leads to:
@@
Our values of $\Gamma$ resemble the highest among those recently measured in NLSy1 galaxies by eROSITA (Grünwald et al., 2023). Maiolino et al. (2024) correctly suggested that an intrinsic X-ray weakness in the LRDs could be caused by a steep, extremely X-ray soft SED, similar to that of NLSy1. Our study confirms that this is the case. Care needs to be taken when choosing the photon index $\Gamma$ to characterize the (restframe) bolometric corrections of the LRDs. Choosing too low values, especially if $\Gamma<2$, will lead to overestimating the restframe bolometric correction. Values in the range $\Gamma=1.7-1.9$, which are typical of sub-Eddington accretion, have been thus far used in the literature to characterize the X-ray weakness of the LRDs (Yue et al., 2024; Maiolino et al., 2024; Ananna et al., 2024).
It is important to note that the spectral descriptors, i.e., $k_{X}$, $\,{\alpha_{\rm ox}}$, and $\Gamma$, derived from our GRRMHD simulations are compatible with those recently observed in super-Eddington accreting SMBHs, both in the local and in the high-$z$ Universe. For example, Laurenti et al. (2022) studied a sample of low-$z$ AGN ($0.4\leq z\leq 0.75$) with bolometric luminosities of $\sim 10^{46}\rm\,erg\,s^{-1}$ and selected for accretion levels close to Eddington, with $0.9\leq\,{f_{\rm Edd}}\leq 1.2$; they found substantial values of the bolometric correction, ranging from $k_{X}=50$ to $k_{X}=5100$, and with a median value of $k_{X}=175$. They observed, however, smaller values of the photon index, i.e., $\Gamma\sim 2$, in agreement with the values found by another study of local, super-Eddington accreting black holes (Maithil et al., 2024). However, the values of $\Gamma$ for super-Eddington accreting SMBHs are higher than those for sub-Eddington ones. In the high-$z$ Universe, Zappacosta et al. (2023) selected highly-accreting AGN at $z>6$ in the HYPERION sample, finding steep values of the photon index $\Gamma=2.4\pm 0.1$, and up to a value of $\Gamma=3.03^{+1.08}_{-0.89}$, which is compatible with the median value for our SEDs.
@@ 4.4 The Cause of X-ray Weakness: Mildly Super-Eddington, Low Spin, and Beaming
Despite a similar Eddington ratio, the SEDs are markedly different. In particular, the X-ray strong SED is characterized by a much harder X-ray spectrum, with a significant bump extending at (restframe) frequencies of $10^{19}-10^{20}$ Hz for viewing angles close to the pole (i.e., $<30^{\circ}$). This high-energy bump is caused by Compton scattering and Lorentz boosting the photons originating from the hottest parts of the accretion flow, as shown in Sec. 4.5. Even the SED averaged over the solid angle (shown in red) is hard in the X-ray. On the contrary, the X-ray weak SED does not show any high-energy bump, except, minimally, for a viewing angle of $10^{\circ}$. Hence, the emission is much softer in the X-rays. The principal difference between the two models is the black hole spin. The $a=0.9$ model has a powerful jet and correspondingly has a large X-ray luminosity. The $a=0$ model has a much weaker jet and is much more X-ray-quiet.
Figure 4: Top Row: Comparison (at $z=0$) between typical X-ray weak (left, mildly super-Eddington with zero spin) and X-ray strong (right, mildly super-Eddington with high spin) SEDs. Bottom Row: Same as in the top row, but at $z=6$. The 8 gray lines represent the spectrum for lines of sight ranging from $80^{\circ}$ (bottom) to $10^{\circ}$ (top) from the pole. The red line indicates the spectrum averaged over $4\pi$ steradians. In the bottom panel, the arrows indicate typical flux limits for deep-field X-ray surveys with Chandra ($2-10$ keV) and AXIS ($0.2-10$ keV). Some of these SEDs extend up to $\sim 1$ MeV.
@@
The situation is radically different for a source characterized by an X-ray weak SED. Chandra can barely detect only sources observed near the pole, and it is $\sim 2$ orders of magnitude in flux away from detecting the mean SED. At the flux limit of $10^{-17}\,\rm erg\,s^{-1}\,cm^{-2}$, the mean SED is too soft, rendering these sources impossible to detect with Chandra.
@@
Next-generation X-ray observatories, such as AXIS (Reynolds et al., 2023; Cappelluti et al., 2024), can detect such weak sources. Assuming an AXIS deep field with a flux limit of $10^{-18}\,\rm erg\,s^{-1}\,cm^{-2}$ (Marchesi et al., 2020; Cappelluti et al., 2024), and an energy range of $0.2-10$ keV, AXIS will detect the mean SED of a typical X-ray weak source at $z\sim 6$ easily and, most likely, from any viewing angle. Note that the SEDs shown in Fig. 4 belong to sources that are easily detectable by JWST/NIRSpec, assuming a flux limit of $\approx 5.5\times 10^{-19}\,\rm erg\,s^{-1}cm^{-2}$ for a $3\sigma$ detection at $\sim 1\,\mu m\approx 3\times 10^{14}\,\rm Hz$ (Pacucci et al., 2023).
@@
Extending the analysis of the SED to lower frequencies, we note that recent studies (e.g., Mazzolari et al. 2024) point out that LRDs that are X-ray weak are also radio-weak. Regarding the predicted radio luminosity, our simulations do not account for non-thermal jet emission, as both KORAL and HEROIC assume thermal gas and focus on thermal synchrotron and Bremsstrahlung radiation. Nevertheless, assuming that the radio emission of the jet is proportional to the jet power, $P_{\rm jet}$, we estimate that the X-ray strong model, with $P_{\rm jet}\approx 0.9\dot{M}c^{2}$, would be more radio luminous than the X-ray weak model, which has $P_{\rm jet}\approx 0.08\dot{M}c^{2}$.
@@ 4.5 X-ray Weak and Strong Accretion Structures
Figure 6: Comparison between accretion structures in the poloidal plane (the black hole is at $(0,0)$ and the jets are oriented in the vertical direction) for our reference cases of X-ray weak (left) and X-ray strong (right) SEDs (see Sec. 4.4). The rows are, from top to bottom: temperature, gas density, scattering optical depth, and $\beta\gamma$. X-ray strong accretion structures are significantly hotter and more relativistic.
@@
In general, for sub-Eddington accretion rates, the disk is thinner (Novikov & Thorne, 1973; Shakura & Sunyaev, 1973), and a more significant fraction of the hot inner region is visible, leading to a more substantial X-ray component in the observed SED, especially if there is an optically thin hot corona above the disk. At super-Eddington accretion rates, although the total luminosity increases, the X-ray contribution diminishes because the hot regions are more obscured by the thicker disk (in Fig. 6, notice that extreme values of the optical depth occupy a larger fraction of the solid angle in the X-ray weak case). This angle dependence implies that the apparent fraction of AGNs with intrinsic strong X-ray emission may be underestimated, as many may be viewed at angles where X-rays are suppressed. This is consistent with the explanation of the X-ray weakness provided by Maiolino et al. (2024), which requires significant covering factors. Additionally, the third row of panels in Fig. 6 shows the scattering optical depth out to $z=100r_{g}$, illustrating the transverse optical depth (integrated over $\theta$ from the pole at a constant radius). These results indicate that X-rays cannot escape laterally, suggesting significant opacity in the transverse direction of the jet.
The main differences in our super-Eddington simulations are shown in the temperature field (top row) and the $\beta\gamma$ field (bottom row), where $\beta=v/c$ and $\gamma$ is the Lorentz factor. The X-ray strong accretion structures display significantly higher values for the temperature and $\beta\gamma$, indicating that the jet is relativistic and more powerful. The combination of these two effects leads to enhanced X-ray emission. Higher spin values generally lead to more efficient jet production, as the spin influences the inner accretion disk’s structure and energy extraction efficiency (see Ricarte et al. 2023 for detailed results on jet efficiencies as a function of black hole spin for super-Eddington accretion).
The temperature distribution within the accretion disk is particularly revealing: the hottest regions, with $T>10^{8}\,\rm K$, are located at the base of the jet and on the wall of the funnel and are the primary sources of X-ray emission. Hence, observers aligned with the jet would predominantly see high-energy X-rays. Conversely, observers at larger inclination angles are screened from the hottest gas by the geometrically thick disk and see cooler regions, resulting in spectra dominated by UV and optical emissions.
In our GRRMHD simulations, the X-rays come from the base of the jet and the wall of the funnel, which is consistent with the lamppost interpretation of the X-ray “coronae" in AGN (see, e.g., Matt et al. 1991; Martocchia & Matt 1996; Miniutti & Fabian 2004). Our simulations (similar to other studies using GRRMHD simulations) do not include an explicit treatment of the X-ray reflection component from an extended corona in the polar region.
@@ 4.6 Explaining the X-ray Weakness of the LRDs
We can now apply our super-Eddington SEDs to the case of the LRDs. Figure 7 shows the bolometric correction $k_{X}^{z=6}$ in the observed $2-10$ keV frame for our sample of super-Eddington SEDs, calculated at $z=6$. Note that the bolometric luminosities of our SEDs are compatible with those observed. We display the entire sample of $12$ simulations $\times$ $8$ viewing angles, totaling $96$ values of the bolometric correction. These are compared to (i) lower limits on the bolometric corrections for LRDs at $114<z<$ 11 derived from Maiolino et al. (2024) and Yue et al. (2024); (ii) the X-ray stacking of the Maiolino et al. (2024) sample, which provides the most robust constraint; (iii) the standard (restframe) relation between the bolometric luminosity and the $2-10$ keV bolometric correction presented in Duras et al. (2020); (iv) a density distribution summarizing the (restframe) location in the $L_{\mathrm{bol}}-k_{X}$ plane of standard, low-$z$ Type 1 AGN from Lusso et al. (2020). Note that the data points from Maiolino et al. (2024) and Yue et al. (2024) have been shifted to the observed $2-10$ keV range to be comparable to our SEDs, using the redshift of each source, and $\Gamma=1.7$ and $\Gamma=1.8$, respectively, as indicated in their study.
@@
Figure 7: In the plane $\log_{10}L_{\mathrm{bol}}$ vs. X-ray bolometric correction (in the observed $2-10$ keV range, for $z=6$), we compare the X-ray weakness of our super-Eddington SEDs with that of high-$z$ LRD populations, and some well-known low-$z$ samples. Teal dots represent the bolometric corrections derived in this study from super-Eddington SEDs at $z=6$ (i.e., the median redshift from LRDs catalogs). Note that the bolometric luminosities of our SEDs are compatible with those observed. Orange circles indicate LRD populations at $114<z<$ 11 from Maiolino et al. (2024) and Yue et al. (2024), with arrows showing lower limits, while the red star marks the stacked data point (for sources at $z>4$) from Maiolino et al. (2024), which provides the most robust constraint. The dark orange circle displays the location of GN-28074, which is the most X-ray weak among all AGN discovered by JWST ($z=2.26$, Juodžbalis et al. 2024a). Note that each data point has been shifted to the observed $2-10$ keV range to be comparable to our SEDs. The solid black line (with its spread) represents the restframe $k_{X}(L_{\mathrm{bol}})$ relation from Duras et al. (2020). The grey distribution in the background shows low-$z$ Type 1 AGN data from Lusso et al. (2020).
@@
Remarkably, our models predict that the restframe optical and UV spectra would still exhibit BLRs, potentially with higher velocities and implied line widths due to strong winds from the super-Eddington accreting SMBH. There is recent observational evidence of strong winds launched from super-Eddington accreting AGN (Vietri et al., 2020, 2022; Ding et al., 2022), supported by robust investigations on the theoretical side (Yang et al., 2014; Sądowski et al., 2015; Sądowski, 2016; Yang et al., 2023; Zhang et al., 2024). This implies that BLRs could appear artificially broadened by these winds, leading to an overestimation of the black hole mass and an underestimation of the Eddington ratio.